Circular No. 8605

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2005en AND 2005eo IN UGC 4132 Two apparent supernova in UGC 4132 have been discovered on un-

filtered CCD images: both objects independently by T. Puckett and M. Peoples (cf. IAUC 8601) and by the Lick Observatory Supernova Search program (KAIT, via R. R. Prasad and W. Li, with 2005eo first spotted by E. Lee on the Sept. 29.54 image; cf. IAUC 8604). SN 2005eo was also independently found by P. Berlind while undertaking spectroscopy for 2005en on Sept. 29 at Mt. Hopkins. The tabulated data below are from Puckett.

SN  2005 UT  α_{2000}  δ_{2000}  Mag.  Offset
2005en Sept. 27.19 7^{h}59^{m}12.^{s}61 +32^{o}55′00′′.5 17.5 5′′.8 W, 6′′.9 N
2005eo Sept. 27.19 7 59 13.94 +32 55 19.7 18.3 11′′.0 E, 26′′.1 N

Li provides position end figures 12^{s}.64, 54′59′′.4 for SN 2005en, and 13^{s}.98, 18′′.4 for 2005eo. Approximate magnitudes for 2005en: 2005 Mar. 2 UT, [20.0 (Puckett); 25, [20.0 (Puckett); Apr. 4.15, [19.5 (Li); Sept. 28.51, 17.0 (Li). Approximate magnitudes for 2005eo: Mar. 2, [20.0 (Puckett); 25, [20.0 (Puckett); Apr. 4.15, [19.5 (Li); Sept. 28.51, 17.5 (Li); 29.17, 18.0 (S. Faworski, Elizabeth, IL, via Puckett); 29.54, 17.5 (Li).

M. Modjaz, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics; and T. Matheson, National Optical Astronomy Observatory, report that a spectrogram (range 350–740 nm) of SN 2005en, obtained by Berlind on Sept. 29.50 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST), reveals it to be a type-II supernova, with Hα and Hβ P-Cgy profiles and other elements in absorption. Adopting the NED recession velocity of 5227 km/s (from Falco et al. 1999, PASP 111, 438 ) for the host galaxy, the expansion velocity derived from the minimum of the Hβ line is 7000 km/s. A spectrogram of 2005eo, taken by Berlind on Sept. 29.55, reveals it to be a type-Ic supernova around maximum; the spectrum is similar to spectra of SN 1994I (Filippenko et al. 1995, Ap.J. 450, L11) taken around maximum.

COMETS 171P/SPAHR AND 172P/YEUNG Permanent numberings have been given recently (cf. MPC 54797) to

two additional comets: 171P/1998 W1 = 2005 R3 (cf. IAUC 8599) and 172P/2001 CB_{40} = 2002 BV (cf. IAUC 7896).

2005 September 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green